OpenStudy (anonymous):

helpp pleasee :") A fair cube with faces 6, 4, 3, 7, 9, and 2 is rolled four times. What is the probability that all rolls will result in an even number?

OpenStudy (kropot72):

This can be modelled by the binomial distribution. Do you understand the binomial distribution?

OpenStudy (anonymous):

no

OpenStudy (kropot72):

There are 3 even numbers out of the 6 total numbers. The probability of an even number is 3/6 = 0.5.\[P(4\ even\ numbers)=\left(\begin{matrix}4 \\ 4\end{matrix}\right)0.5^{4}0.5^{0}=0.5^{4}\]

OpenStudy (anonymous):

i got it it

OpenStudy (kropot72):

Good! A simpler way of looking at the problem is to consider that the probability of an even number on each roll is 0.5. There are 4 rolls each with a probability of 0.5 so the probability of the result of an even number on all 4 rolls is 0.5 * 0.5 * 0.5 * 0.5 =\[0.5^{4}\]